(2+√y)^2-(2-√y)^2=8√y but in your solution it is equal to 4√y what's the reason.


1
Expert's answer
2021-08-05T13:50:22-0400

"(2+\sqrt{y})^2-(2-\sqrt{y})^2\\
=((2+\sqrt{y})+(2-\sqrt{y}))((2+\sqrt{y})-(2-\sqrt{y}))\\
=(2+2+\sqrt{y}-\sqrt{y})(2-2+\sqrt{y}+\sqrt{y})\\
=(4)(2\sqrt{y})\\
=8\sqrt{y}\\
\hspace{0.0cm}\\
\textsf{Hence,}\\
(2+\sqrt{y})^2-(2-\sqrt{y})^2=8\sqrt{y}\not=4\sqrt{y}\\"